Circular No. 7150

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

POSSIBLE SUPERNOVA IN NGC 3198 P. Garnavich, S. Jha, and R. Kirshner, Harvard-Smithsonian Center

for Astrophysics, report that spectra of the possible supernova in NGC 3198 (cf. IAUC 7149) were obtained with the 1.5-m Tillinghast telescope by M. Calkins on Apr. 21.2 UT. The spectrum is not clearly that of a supernova. The continuum peaks broadly around 500 nm and declines more quickly in the blue than toward the red. Balmer emission lines of Hα, Hβ, and Hγ are visible. The Hα line has a narrow component sitting on a broader base, with a full width at the continuum of 2000 km/s. CCD images obtained by W. Brown with the Fred L. Whipple Observatory 1.2-m telescope (at the same time as the spectra) give V = 18.5. The object has not risen in brightness significantly over the six days since discovery (IAUC 7149), as would be expected from a supernova caught at an early stage.

A. V. Filippenko and T. Matheson, University of California at Berkeley, report: “Inspection of a CCD spectrum (range 400–860 nm) of the possible supernova in NGC 3198, obtained on Apr. 21 by P. Guhathakurta and A. Szomoru (University of California at Santa Cruz) with the 3-m Shane reflector at Lick Observatory, reveals Balmer emission lines near the galaxy redshift (630 km/s). The lines are unresolved, or perhaps barely resolved (FWHM = 1000 km/s), in the low-resolution data. If this object is a supernova, it is very subluminous: absolute magnitude ∼ −12 at an assumed distance of 9 Mpc. This is much more powerful than an ordinary nova, but within the range of some kinds of luminous blue variable stars. The exact nature of the object therefore remains ambiguous.”

SUPERNOVA 1999bv IN MGC +10-25-14 S. Jha, P. Garnavich, P. Challis, and R. Kirshner, Center for Astro-

physics, report that a spectrum of SN 1999bv, obtained by M. Calkins on Apr. 21.5 UT with the F. L. Whipple Observatory (FLWO) 1.5-m telescope (+ FAST spectrograph), indicates that it is probably a type-Ib/c supernova. The spectrum exhibits a flat continuum with broad (FWHM 10 000 km/s) emission features observed at 667, 600, 475, and 402 nm. This is generally similar to a spectrum of the type-Ib/c SN 1988L when it was presumed to be a few weeks after maximum light (Filippenko 1988, A.J. 96, 1941). Narrow Hα emission from the host galaxy yields a recession velocity of 5510 km/s. Concurrent images taken by W. Brown with the FLWO 1.2-m telescope (+ 4shooter) yield V = 18.6 for the supernova. SN 1999bv is within the 2.5σ error box of GRB 990302 (BATSE Trigger 7440).

1999 April 22  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green